# Optimal control of the fidelity coefficient in a Cahn–Hilliard image inpainting model

### v1

@@ Abstract
We consider an inpainting model proposed by A. Bertozzi et al. which is based on a Cahn–Hilliard-type equation. This equation describes the evolution of an order parameter $u\in[0,1]$ representing an approximation of the original image which occupies a bounded two-dimensional domain $\Omega$. The given image $g$ is assumed to be damaged in a fixed subdomain $D\subset\Omega$ and the equation is characterized by a linear reaction term of the form $\lambda(u-g)$. Here $\lambda=\lambda_{0}\chi_{\Omega\setminus D}$ is the so-called fidelity coefficient, $\lambda_{0}$ being a strictly positive bounded function. The idea is that, given an initial image $u_{0}$, $u$ evolves towards $g$ and this process properly diffuses through the boundary of $D$ restoring the damaged image, provided that $\lambda_{0}$ is large enough. Here, we formulate an optimal control problem based on this fact, namely our cost functional accounts for the magnitude of $\lambda_{0}$. Assuming a singular potential to assure that $u$ takes its values in $[0,1]$, we first analyse the control-to-state operator and prove the existence of at least one optimal control, establishing the validity of first-order optimality conditions. Then, under suitable assumptions, we demonstrate second-order optimality conditions. All these results depend on the existence and uniqueness of a strong solution, which we obtain thanks to the strict separation property from pure phases.
@@ 1 Introduction
One possible drawback of model (1.1)–(1.4) is that, with a polynomial double-well potential as $W(u)=u^{2}(u-1)^{2}$, one cannot guarantee that the solution $u$ takes its values in $[0,1]$. However, this issue can be solved by reformulating the problem using a physically relevant potential, namely, the so-called logarithmic-type double-well (or Flory–Huggins) potential (see [12, 25, 26]). In some situations, potentials of this kind were also shown to provide better inpainting results and faster-converging algorithms (see [12]). For convenience, we now change most of the notation used so far in order to work in the physical interval $[-1,1]$, instead of $[0,1]$, since it is more common in the literature on Cahn–Hilliard-type equations. Clearly, this does not change anything in applications, since one can easily switch between the two formulations by setting $\varphi=2(u-\frac{1}{2})$. More precisely, we replace $W(u)$ by
@@
for a fixed constant $m_{*}\in(0,1)$. Therefore, the inpainted image is now a function $\overline{\varphi}:\Omega\to[-1,1]$, which should always be approximately equal to ±mplus-or-$\pm m_{*}$ up to a diffuse interface of width $\varepsilon$ (see [25, Rem. 2.1]). Indeed, we find $\overline{\varphi}=\varphi(T)$ by solving the following evolutionary problem:
@@ 2 Preliminaries and well-posedness results
$\lVert u\rVert^{2}_{W}:=\lVert u\rVert^{2}_{H}+\lVert\Delta u\rVert^{2}_{H},$
where $\Delta$ denotes the Laplacian. We also need to define the generalised mean value of a function $v$ as
@@
$\langle\mathcal{N}u,v\rangle:=\int_{\Omega}\nabla u\cdot\nabla v\,\mathrm{d}x,$
is positive definite and self-adjoint. In particular, by the Lax–Milgram theorem and Poincaré–Wirtinger’s inequality, the inverse operator $\mathcal{N}^{-1}:V^{*}_{0}\to V_{0}$ is well defined and $u=\mathcal{N}^{-1}g$ if $u$ is the weak solution to the elliptic boundary value problem
$\begin{cases}-\Delta u=g\quad&\hbox{in $\Omega$,}\\ \partial_{\mathbf{n}}u=0\quad&\hbox{on $\partial\Omega$.}\end{cases}$
@@ 2.1 Weak and strong solutions
    $F_{0}^{\prime\prime}(s)\geq c_{0}>0,\quad F_{0}^{\prime\prime}(s)\leq e^{C(1+\lvert F_{0}^{\prime}(s)\rvert)},$ (2.2)
@@
Assumptions A1A5 are the same used in [25] to show the existence of strong solutions to (1.7)–(1.10), as we will recall below. Observe that A2 is satisfied by the logarithmic double-well potential (1.6) introduced above. We also mention that (2.2) can be further weakened in view of the recent results regarding the validity of the separation property in two spatial dimensions (see [21, Thm. 3.3]). Moreover, for A3, we refer again to [25, Rem. 2.1].
We first recall the result contained in [25, Proposition 2.5] on the existence of weak solutions to (1.7)–(1.10) (cf. also [26]).
@@
$\varphi\in H^{1}(0,T;V^{*})\cap L^{\infty}(0,T;V)\cap L^{4}(0,T;W)\cap L^{2}(0,T;W^{2,p}(\Omega)),$
@@
$\lvert\varphi(x,t)\rvert<1\hbox{ a.e.\leavevmode\nobreak\ in $Q_{T}$,}$
@@
$\mu=-\Delta\varphi+F^{\prime}(\varphi)\quad\hbox{a.e.\leavevmode\nobreak\ in $Q_{T}$,}$
@@
Next, we recall the result we will use in the sequel to define the control-to-state operator, namely, the well-posedness of problem (1.7)–(1.10) with respect to strong solutions (cf. [25, Theorem 3.2]). We point out that a key step to obtain higher regularity for solutions to (1.7)–(1.10) is to establish the strict separation property, that is, proving that $\varphi$ stays uniformly away from ±$\pm 1$ in $[0,T]$ (see [21] and references therein). This basic property allows us to handle the singular potential as if it were a globally Lipschitz nonlinearity.
@@
In the original [25, Theorem 3.2] only the instantaneous strict separation property was shown, namely that for any $\sigma>0$ there exists $\delta>0$ such that
@@
However, we observe that our assumptions on $\varphi_{0}$ and $\mu_{0}$ already entail the existence of a $10<\delta_{0}<$ 1 such that $\varphi$ is strictly separated locally in time (see [23, Rem. 3.9] and references therein). Thus the strict separation property holds up to time $t=0$ (see [23, Rem. 3.11]), as shown in (2.3) for some $0<\delta<\delta_{0}$. In particular, we stress that $\delta_{0}$ depends only on the initial energy, therefore it can be chosen uniformly if the energy of the initial guess $\varphi_{0}$ is uniformly bounded.
The analysis conducted in [25] shows, in particular, that the solutions to (1.7)–(1.10) regularise in finite time (see [25, Thm. 3.1]). Hence, for the purposes of our analysis, we can assume to start with an initial guess $\varphi_{0}$ satisfying
@@ 2.2 Continuous dependence on the fidelity coefficient
In order to study the optimal control problem (CP) in the next section, we need to define a control-to-state operator. Along with the existence and uniqueness of a strong solution given by Theorem 2.3 and the validity of the separation property (2.3), we also need to establish a continuous dependence of the (strong) solution itself on the fidelity coefficient $\lambda$. This is given by
@@
We argue as in the proof of [25, Theorem 3.2]. Let us set
@@
$\partial_{t}\varphi-\Delta\mu=\lambda(f-\varphi_{2})-\lambda_{1}\varphi\quad$ a.e. in $Q_{T}$, (2.10)
$\mu=-\Delta\varphi+(F^{\prime}(\varphi_{1})-F^{\prime}(\varphi_{2}))\quad$ a.e. in $Q_{T}$, (2.11)
$\partial_{\mathbf{n}}\varphi=\partial_{\mathbf{n}}\mu=0\quad$ a.e. on $\Sigma_{T}$, (2.12)
@@
Likewise, by comparison in (2.10), one can also easily show that
@@
To estimate the first term on the right-hand side of (2.18), we initially rewrite it better by adding and subtracting some terms and integrating by parts. Then, we employ Hölder and Young’s inequalities, Sobolev embeddings and Remark 2.5 to deduce that
@@
$(\lambda\varphi_{2},\partial_{t}\varphi)_{H}\leq\lVert\lambda\rVert_{H}\lVert\varphi_{2}\rVert_{H}\lVert\partial_{t}\varphi\rVert_{H}\leq\frac{1}{8}\lVert\partial_{t}\varphi\rVert^{2}_{H}+C\lVert\lambda\rVert^{2}_{H}.$
@@ 3 Existence of an optimal control
We mention that all the analysis we will carry out in the following can be extended in a straight-forward way to the more general case in which the last term in the cost functional $\mathcal{J}$ is replaced by
@@
for any $r>0$. Here we just choose $r=2$ for convenience.
@@
In particular, thanks to well-known compact embeddings (see [32, Section 8, Corollary 4]), it follows that $\varphi_{n}\to\overline{\varphi}$ strongly in $C^{0}([0,T];H^{s}(\Omega))$ for any $20<s<$ 2 which implies that $\varphi_{n}\to\overline{\varphi}$ strongly in $C^{0}(\overline{Q_{T}})$, since $H^{s}(\Omega)\hookrightarrow C^{0}(\overline{Q_{T}})$ if $s>1$ in dimension two. Moreover, by uniform convergence, $\overline{\varphi}$ still satisfies the strict separation property, namely there exists $\delta>0$ such that $\lVert\overline{\varphi}\rVert_{C^{0}(\overline{Q_{T}})}\leq 1-\delta$. This fact, by the continuity of $F^{\prime}$ inside $(-1,1)$, implies that
@@ 4 First-order optimality conditions
Then, by Theorem 2.3, we have that
$\mathcal{S}:\mathcal{U}_{\text{ad}}\to\mathbb{X}\quad\hbox{is well defined,}$,
and, by Theorem 2.6, we also have that
$\mathcal{S}:\mathcal{U}_{\text{ad}}\to\mathbb{Y}\quad\hbox{is Lipschitz continuous.}$.
Our next step is to show the Fréchet differentiability of $\mathcal{S}$ in order to differentiate the cost functional and derive the first-order necessary optimality conditions. Indeed, in this section, we first study the linearised system and we prove the Fréchet differentiability of $\mathcal{S}$. Then, we prove the well-posedness of the adjoint system which allows us to show the first-order optimality conditions.
@@ 4.1 Linearised system
As an ansatz for the Fréchet derivative of the control-to-state operator, we now introduce the linearised system corresponding to our control problem. Indeed, we fix an optimal state $(\overline{\varphi},\overline{\mu})$ corresponding to $\overline{\lambda}_{0}\in\mathcal{U}_{\text{ad}}$ and we linearise near $\overline{\lambda}_{0}$:
@@
with $h_{0}\in L^{\infty}(\Omega\setminus D)$ such that $\lVert h_{0}\rVert_{L^{\infty}(\Omega\setminus D)}<\lambda_{\text{min}}$. Then, by approximating the non-linearities at the first order of their Taylor expansion, it is easy to see that $(\xi,\eta)$ formally satisfies the problem
@@
The proof can be carried out rigorously through a Faedo–Galerkin discretisation scheme using finite-dimensional spaces generated by the eigenvectors of the operator $\mathcal{N}$. However, being the system linear, this procedure is quite standard, thus we proceed by providing only formal a priori estimates, leaving the details to the interested reader.
@@
$\leq\lVert F_{1}^{\prime\prime}(\overline{\varphi})\rVert_{L^{\infty}(\Omega)}\lVert\nabla\xi\rVert_{H}\lVert\nabla\eta\rVert_{H}+\lVert F_{1}^{\prime\prime\prime}(\overline{\varphi})\rVert_{L^{\infty}(\Omega)}\lVert\nabla\overline{\varphi}\rVert_{L^{4}(\Omega)}\lVert\nabla\eta\rVert_{H}\lVert\xi\rVert_{H}$
@@
$(h(f-\overline{\varphi}),\partial_{t}\xi)_{H}\leq\lVert h\rVert_{H}\left(\lVert f\rVert_{L^{\infty}(\Omega)}+\lVert\overline{\varphi}\rVert_{L^{\infty}(\Omega)}\right)\lVert\partial_{t}\xi\rVert_{H}\leq\frac{1}{4}\lVert\partial_{t}\xi\rVert^{2}_{H}+C\lVert h\rVert^{2}_{H}$
@@
Then, integrating (4.10) in $(0,t)$, for any $t\in(0,T)$, and using all the above estimates, we infer that
@@
where $g_{1}$ and $g_{2}$ are given sources ($g_{1}=h(f-\overline{\varphi})$ and $g_{2}=0$ in (4.1)–(4.4)). In particular, we the following results can be proven:

### v2

@@ Abstract
We consider an inpainting model proposed by A. Bertozzi et al., which is based on a Cahn–Hilliard-type equation. This equation describes the evolution of an order parameter that represents an approximation of the original image occupying a bounded two-dimensional domain. The given image is assumed to be damaged in a fixed subdomain, and the equation is characterised by a linear reaction term. This term is multiplied by the so-called fidelity coefficient, which is a strictly positive bounded function defined in the undamaged region. The idea is that, given an initial image, the order parameter evolves towards the given image and this process properly diffuses through the boundary of the damaged region, restoring the damaged image, provided that the fidelity coefficient is large enough. Here, we formulate an optimal control problem based on this fact, namely, our cost functional accounts for the magnitude of the fidelity coefficient. Assuming a singular potential to ensure that the order parameter takes its values in between 00 and $1$, we first analyse the control-to-state operator and prove the existence of at least one optimal control, establishing the validity of first-order optimality conditions. Then, under suitable assumptions, we demonstrate second-order optimality conditions.
@@ 1 Introduction
One possible drawback of model (1.1)–(1.4) is that, with a polynomial double-well potential as $W(u)=u^{2}(u-1)^{2}$, one cannot guarantee that the solution $u$ takes its values in $[0,1]$. However, this issue can be solved by reformulating the problem using a physically relevant potential, namely, the so-called logarithmic-type double-well (or Flory–Huggins) potential (see [13, 28, 29]). In some situations, potentials of this kind were also shown to provide better inpainting results and faster-converging algorithms (see [13]). For convenience, we now change most of the notation used so far in order to work in the physical interval $[-1,1]$, instead of $[0,1]$, since it is more common in the literature on Cahn–Hilliard-type equations. We call $\varphi$ this new phase-field variable in $[-1,1]$. Clearly, this does not change anything in applications, since one can easily switch between the two formulations by setting $\varphi=2(u-\frac{1}{2})$. More precisely, we replace $W(u)$ by
@@
for a fixed constant $m_{*}\in(0,1)$. Therefore, the inpainted image is now a function $\overline{\varphi}:\Omega\to[-1,1]$, which should always be approximately equal to ±mplus-or-$\pm m_{*}$ up to a diffuse interface of width $\varepsilon$ (see [28, Rem. 2.1]). Indeed, we find $\overline{\varphi}=\varphi(T)$ by solving the following evolutionary problem:
@@ 2 Preliminaries and well-posedness results
$\lVert w\rVert^{2}_{W}:=\lVert w\rVert^{2}_{H}+\lVert\Delta w\rVert^{2}_{H},$
where $\Delta$ denotes the Laplace operator. We also need to define the generalised mean value of a function $v$ as
@@
$\langle\mathcal{N}w,v\rangle:=\int_{\Omega}\nabla w\cdot\nabla v\,\mathrm{d}x,$
is positive definite and self-adjoint. In particular, by the Lax–Milgram theorem and Poincaré–Wirtinger’s inequality, the inverse operator $\mathcal{N}^{-1}:V^{*}_{0}\to V_{0}$ is well defined and $w=\mathcal{N}^{-1}g$ if $w\in V_{0}$ is the weak solution to the elliptic boundary value problem
$\begin{cases}-\Delta w=g\quad&\hbox{in $\Omega$,}\\ \partial_{\mathbf{n}}w=0\quad&\hbox{on $\partial\Omega$.}\end{cases}$
@@ 2.1 Weak and strong solutions
    $F_{0}^{\prime\prime}(s)\geq c_{0}>0,\quad F_{0}^{\prime\prime}(s)\leq e^{C(1+\lvert F_{0}^{\prime}(s)\rvert)}.$ (2.2)
@@
Assumptions A1A5 are the same used in [28] to show the existence of strong solutions to (1.7)–(1.10), as we will recall below. Observe that A2 is satisfied by the logarithmic double-well potential (1.6) introduced above. We also mention that (2.2) can be further weakened in view of the recent results regarding the validity of the separation property in two spatial dimensions (see [24, Thm. 3.3]). Moreover, for A3, we refer again to [28, Rem. 2.1].
We first recall the result contained in [28, Proposition 2.5] on the existence of weak solutions to (1.7)–(1.10) (cf. also [29]).
@@
$\varphi\in H^{1}(0,T;V^{*})\cap L^{\infty}(0,T;V)\cap L^{4}(0,T;W)\cap L^{2}(0,T;W^{2,p}(\Omega))\quad\hbox{for any $p\in[2,+\infty)$,}$
@@
$\lvert\varphi(x,t)\rvert<1\quad\hbox{for a.e.\leavevmode\nobreak\ $(x,t)\in Q_{T}$,}\quad\lvert\varphi_{\Omega}(t)\rvert<1\quad\hbox{for any $t\in(0,T)$,}$
@@
$\mu=-\varepsilon\Delta\varphi+\frac{1}{\varepsilon}F^{\prime}(\varphi)\quad\hbox{a.e.\leavevmode\nobreak\ in $Q_{T}$,}$
@@
Next, we recall the result we will use in the sequel to define the control-to-state operator, namely, the well-posedness of problem (1.7)–(1.10) with respect to strong solutions (cf. [28, Theorem 3.2]). We point out that a key step to obtain higher regularity for solutions to (1.7)–(1.10) is to establish the strict separation property, that is, proving that $\varphi$ stays uniformly away from ±$\pm 1$ in $[0,T]$ (see [24] and references therein). This basic property allows us to handle the singular potential as if it were a globally Lipschitz nonlinearity.
@@
In the original [28, Theorem 3.2] only the instantaneous strict separation property was shown, namely that for any $\sigma>0$ there exists $\delta>0$ such that
@@
However, we observe that our assumptions on $\varphi_{0}$ and $\mu_{0}$ already entail the existence of a $10<\delta_{0}<$ 1 such that $\varphi$ is strictly separated locally in time (see [26, Rem. 3.9] and references therein). Thus the strict separation property holds up to time $t=0$ (see [26, Rem. 3.11]), as shown in (2.3) for some $0<\delta<\delta_{0}$. In particular, we stress that $\delta_{0}$ depends only on the initial energy, therefore it can be chosen uniformly if the energy of the initial guess $\varphi_{0}$ is uniformly bounded.
The analysis conducted in [28] shows, in particular, that the solutions to (1.7)–(1.10) regularise in finite time (see [28, Thm. 3.1]). Hence, for the purposes of our analysis, we can assume to start with an initial guess $\varphi_{0}$ satisfying
@@ 2.2 Continuous dependence on the fidelity coefficient
In order to study the optimal control problem (CP) in the next section, we need to define a control-to-state operator. Along with the existence and uniqueness of a strong solution given by Theorem 2.3 and the validity of the separation property (2.3), we also need to establish a continuous dependence of the (strong) solution itself on the fidelity coefficient $\lambda$.
@@
We argue as in the proof of [28, Theorem 3.2]. Let us set
@@
$\partial_{t}\varphi-\Delta\mu=\lambda(f-\varphi_{2})-\lambda_{1}\varphi\quad$ in $Q_{T}$, (2.10)
$\mu=-\Delta\varphi+(F^{\prime}(\varphi_{1})-F^{\prime}(\varphi_{2}))\quad$ in $Q_{T}$, (2.11)
$\partial_{\mathbf{n}}\varphi=\partial_{\mathbf{n}}\mu=0\quad$ on $\Sigma_{T}$, (2.12)
@@
Likewise, by (2.10), one can also easily show that
@@
To estimate the first term on the right-hand side of (2.18), we initially rewrite it by integrating by parts and by adding and subtracting some terms (see the first two equalities below). We stress that the integration by parts does not produce boundary terms, since the separation property and (2.7) ensure that $\partial_{\mathbf{n}}F^{\prime}(\varphi_{i})=F^{\prime\prime}(\varphi_{i})\partial_{n}\varphi_{i}=0$, for $i=1,2$. Then, we employ Hölder and Young’s inequalities, Sobolev embeddings and Remark 2.5 to deduce that
@@
$(\lambda\varphi_{2},\partial_{t}\varphi)_{H}\leq\lVert\lambda\rVert_{H}\lVert\varphi_{2}\rVert_{L^{\infty}(\Omega)}\lVert\partial_{t}\varphi\rVert_{H}\leq\frac{1}{8}\lVert\partial_{t}\varphi\rVert^{2}_{H}+C\lVert\lambda\rVert^{2}_{H}.$
@@ 3 Existence of an optimal control
We mention that all the analysis we will carry out in the following can be extended in a straight-forward way to the more general case in which the last term in the cost functional $\mathcal{J}$ is replaced by the integral of a $C^{2}$ convex function of $\lambda_{0}$. For instance, to penalise large values of $\lambda_{0}$, one can choose
@@
for any $r>0$. Here we just use $r=2$ for convenience.
@@
In particular, thanks to well-known compact embeddings (see [35, Section 8, Corollary 4]), it follows that $\varphi_{n}\to\overline{\varphi}$ strongly in $C^{0}([0,T];H^{s}(\Omega))$ for any $20<s<$ 2 which implies that $\varphi_{n}\to\overline{\varphi}$ strongly in $C^{0}(\overline{Q_{T}})$, since $H^{s}(\Omega)\hookrightarrow C^{0}(\overline{Q_{T}})$ if $s>1$ in dimension two. Moreover, by uniform convergence, $\overline{\varphi}$ still satisfies the strict separation property, namely there exists $\delta>0$ such that $\lVert\overline{\varphi}\rVert_{C^{0}(\overline{Q_{T}})}\leq 1-\delta$. This fact, by the continuity of $F^{\prime}$ inside $(-1,1)$, implies that
@@ 4 First-order optimality conditions
Then, by Theorem 2.3, we have that $\mathcal{S}:\mathcal{U}_{\text{ad}}\to\mathbb{X}$ is well defined and that, by Theorem 2.6, $\mathcal{S}:\mathcal{U}_{\text{ad}}\to\mathbb{Y}$ is Lipschitz continuous with respect to the $H$-norm in $\mathcal{U}_{\text{ad}}$ and the natural one in $\mathbb{Y}$. Our next step is to show the Fréchet differentiability of $\mathcal{S}$ in order to differentiate the cost functional and derive the first-order necessary optimality conditions. Indeed, in this section, we first study the linearised system and we prove the Fréchet differentiability of $\mathcal{S}$. Then, we prove the well-posedness of the adjoint system which allows us to show the first-order optimality conditions.
@@ 4.1 Linearised system
As an ansatz for the Fréchet derivative of the control-to-state operator, we now introduce the linearised system corresponding to our control problem. Indeed, to formulate the linearised system, we fix an optimal control $\overline{\lambda}_{0}\in\mathcal{U}_{\text{ad}}$, corresponding to an optimal state $(\overline{\varphi},\overline{\mu})\in\mathbb{X}$, and we linearise near $\overline{\lambda}_{0}$:
@@
with $h_{0}\in L^{\infty}(\Omega\setminus D)$ such that $\lVert h_{0}\rVert_{L^{\infty}(\Omega\setminus D)}<\lambda_{\text{min}}$ and $\overline{\lambda}_{0}+h_{0}\in\mathcal{U}_{\text{ad}}$. Then, by approximating the non-linearities at the first order of their Taylor expansion, it is easy to see that $(\xi,\eta)$ formally satisfies the problem
@@
The proof can be carried out rigorously through a Faedo–Galerkin discretisation scheme using finite-dimensional spaces generated by the eigenvectors of the operator $\mathcal{N}$. However, being the system linear, this procedure is quite standard, thus we proceed by providing only formal a priori estimates, leaving the details to the interested reader. In particular, due to the linearity of the system, the discrete solutions are $C^{1}$ with respect to time with values in the discrete eigenspaces, thus all the upcoming a priori estimates make sense within the discretisation scheme.
@@
$\leq\lVert F_{1}^{\prime\prime}(\overline{\varphi})\rVert_{L^{\infty}(\Omega)}\lVert\nabla\xi\rVert_{H}\lVert\nabla\eta\rVert_{H}+\lVert F_{1}^{\prime\prime\prime}(\overline{\varphi})\rVert_{L^{\infty}(\Omega)}\lVert\nabla\overline{\varphi}\rVert_{L^{4}(\Omega)}\lVert\nabla\eta\rVert_{H}\lVert\xi\rVert_{L^{4}(\Omega)}$
@@
$(h(f-\overline{\varphi}),\partial_{t}\xi)_{H}\leq\lVert h\rVert_{H}\left(\lVert f\rVert_{L^{\infty}(\Omega)}+\lVert\overline{\varphi}\rVert_{L^{\infty}(\Omega)}\right)\lVert\partial_{t}\xi\rVert_{H}\leq\frac{1}{4}\lVert\partial_{t}\xi\rVert^{2}_{H}+C\lVert h\rVert^{2}_{H},$
@@
Then, integrating (4.10) over $(0,t)$, for any $t\in(0,T)$, and using all the above estimates, we infer that
@@
where $g_{1}$ and $g_{2}$ are given sources ($g_{1}=h(f-\overline{\varphi})$ and $g_{2}=0$ in (4.1)–(4.4)). In particular, the following results can be proven: